OpenStudy (anonymous):

-15+6(-3b+4)+9(2b-1)

Directrix (directrix):

0

OpenStudy (campbell_st):

take it bit by bit.. 6(-3b + 4) = -18b + 24 9(2b -1) = 18b - 9 combining -15 -18b +14 +18b - 9 collect like terms.... I'll leave that for you

OpenStudy (campbell_st):

oops 24 not 14

Directrix (directrix):

-15+6(-3b+4)+9(2b-1) = -15 -18b + 24 + 18b -9 = 0